# Similarity to contraction semigroups:

### v1

@@ 1.1. State of the art, motivations and illustrations
Despite Pisier’s negative result, the study of operators similar to contractions and their sub-classes became a well-developed area, and it continued to bear fruit, see e.g. [5], [16], [22], [32], [59] and [73] as samples. It is instructive to observe that if $T\in\mathcal{L}(H)$ satisfies the so-called Ritt’s resolvent condition $\|(z-T)^{-1}\|\leq C|z-1|^{-1}$ for $z$ outside the unit disc, then the polynomial boundedness of $T$ does imply its similarity to a contraction, [49, Theorem 5.1]. However, there are Ritt operators which are not polynomially bounded, and thus not similar to a contraction, as for instance [49, Proposition 5.2] shows. Among other results linking similarity to contraction to resolvent estimates, polynomial and power boundedness, we note [11], [16], [22] and [31]. However, it is worth noting that polynomial boundedness is a rather strong assumption, and it is often difficult to verify in concrete applications.
@@
Unfortunately, despite its importance for applications, $\mathcal{SC}(H)$ is not well-understood and in contrast to the discrete case there are very few conditions ensuring similarity to a contraction semigroup. At the same time, several pertinent examples were found illustrating the difficulty of the problem. Relying on ideas of Foguel [30], Packel constructed in [65] an example of a bounded $C_{0}$-semigroup $\mathcal{T}\not\in\mathcal{SC}(H),$ with the generator being rather implicit. In [9], Benchimol provided a similar construction of $\mathcal{T}\not\in\mathcal{SC}(H).$ However, he first constructed the generator of $\mathcal{T}$ as a bounded perturbation of a generator of a contraction semigroup, thus making both the generator and the semigroup transparent and explicit. Moreover, among other instructive examples, Chernoff showed in [17, p. 254] how to produce an exponentially stable $C_{0}$-semigroup that does not belong to $\mathcal{SC}(H).$ This somewhat discouraged research on $\mathcal{SC}(H),$ and much later even stronger counterexamples emerged. In particular, Le Merdy observed in [49, Theorem 1.1] that a similar example can be realised for exponentially stable holomorphic semigroup. In addition, in contrast to the case of holomorphic semigroups, one can find $\mathcal{T}\not\in\mathcal{SC}(H)$ whose negative generator is invertible and has bounded imaginary powers. The strongest known example of $\mathcal{T}\not\in\mathcal{SC}(H)$ was again found by Le Merdy in [50], see [36, Section 9.1] for further discussion. In his example, $\mathcal{T}=(T(t))_{t\geq 0}$ is exponentially stable, consists of compact operators for $t>0,$ and is holomorphic of angle $\pi/2.$ (The latter was not stated in [50], but can be easily deduced using e.g. [36, Lemma 9.1.2].) Very recently, among other improvements to Le Merdy’s example, it was shown in [63, Theorem 1.3] that one can construct an even stronger example by replacing exponential decay with nilpotency while preserving compactness, though necessarily omitting holomorphicity. However, holomorphicity is present in another example from [63, Theorem 1.3], where nilpotency is relaxed to quasi-nilpotency.
@@
Among very few positive results, Le Merdy proved in [49, Theorem 1.1] that a sectorially bounded holomorphic semigroup belongs to $\mathcal{SC}(H)$ if and only if its negative generator has bounded imaginary powers, or equivalently admits a bounded $H^{\infty}$-calculus on an appropriate sector. This allowed one to formulate similarity criteria in terms of square function estimates, [36, Theorem 7.3.1], [49, Theorem 4.3], and, a posteriori, invoking resolvent of the generator [13] (where boundedness of imaginary powers was described via integral resolvent bounds). Moreover, aiming at an analogue of Nagy’s theorem, it was shown in [79] that every bounded, uniformly continuous and quasi-compact semigroup in a Hilbert space is similar to a contraction semigroup. Furthermore, [3, Theorem 4.3] showed that the assumption of uniform continuity in [79] can be relaxed to the weaker conditions of analyticity and quasi-contractivity.
@@
The examples by Chernoff and Le Merdy, along with their improvements in [63, Theorem 1.3], demonstrate two key points. First, they show that continuous analogs of the classical criteria by Nagy and Rota fail in a dramatic way. Second, they clarify that our study concerns in fact the joint similarity of the semigroup $(T(t))_{t\geq 0}$ to a family of contractions on $H$, i.e. $\|RT(t)R^{-1}\|\leq 1$ for an invertible $R\in\mathcal{L}(H)$ and $t\geq 0$. This joint similarity does not, in general, follow from the similarity of each individual $T(t),$ $t\geq 0,$ to contraction - unlike the discrete case of the semigroup $(T^{n})_{n\geq 0}.$ As it will be made clear in Section 6 below, the property $\mathcal{T}\in\mathcal{SC}(H)$ depends on the behavior of the similarity constants $C(T(t))$ near zero. For an operator $T\in\mathcal{L}(H)$ its similarity constant is defined as
@@
Quasi-contraction semigroups originate from the classical work [55] and frequently appear in applications. For instance, they play a crucial role in proving the well-posedness of concrete linear PDEs via the (shifted) Lumer-Phillips theorem. They also arise in various stability problems, such as those related to Chernoff’s formula and its applications in numerical analysis, as well as in the study of nonlinear PDEs and their linearizations. To illustrate the concept of quasi-contractivity, observe that every $C_{0}$-semigroup $\mathcal{T}$ with a bounded generator $E$ is quasi-contractive as it satisfies the bound $\|T(t)\|\leq e^{\|E\|t}$ for all $t\geq 0.$ Furthermore, since any bounded perturbation of the generator of a quasi-contraction semigroup results in another quasi-contraction semigroup, bounded perturbations of normal operators whose spectrum lies in some left half-plane generate quasi-contraction semigroups. By a similar argument, one can show that every $C_{0}$-group on $H$ is similar to a quasi-contraction semigroup (cf. [36, Theorem 7.2.8 and Corollary 7.2.9]).
Thus, quasi-contractivity is a significantly weaker requirement than contractivity and naturally arises in semigroup theory. On the other hand, Chernoff’s example mentioned above ([17, p. 254]) actually yields a semigroup not similar to a quasi-contraction semigroup. Furthermore, by [36, Theorem 7.3.13] one can construct an analogous example where, in addition, negative generator is invertible and has bounded imaginary powers. Moreover, as we show in Section 7, the Packel and Le Merdy semigroups do not belong to $\mathcal{SQC}(H)$. Even stronger examples of semigroups outside $\mathcal{SQC}(H)$ can be found in [63]. At the same time, Benchimol’s semigroup is similar to a quasi-contraction semigroup since its generator arises as a bounded perturbation of the generator of a contraction semigroup. Thus the relations between $\mathcal{SC}(H)$ and $\mathcal{SQC}(H)$ appear to be rather intricate, and the aim to clarify them underlines the specifics and novelty of our approach. The paper links quasi-contractivity to the study of semigroups in $\mathcal{SC}(H)$ and sheds new light on their structure, revealing the different roles of small and large times and their interplay. It aims to emphasize these roles and to make the subject accessible to different audiences.
@@ 1.2. Ideas and results
The result can be proved by following the arguments from the proof of its discrete analogue, established in [42]. See also [71, Proposition 4.2] for a more general setting. This characterization, together with its generalization in the spirit of (1.1), is fundamental to our approach. Accordingly, we provide both results with new, more explicit proofs, placing them within the context of representations of (algebraic) semigroups. In particular, we make quotient norms explicit and thereby avoid the indirect argument requiring verification of the parallelogram law, as in [43]. In this way, given $\tau>0$, we infer that $(T(t))_{t\geq\tau}$ is jointly similar to contractions if (1.3) holds only for $t\geq\tau$. This observation becomes crucial in the next step of our program, where we describe $\mathcal{T}=(T(t))_{t\geq 0}$ with at least one $T(t)$ similar to a contraction.
@@
Remark that if $\mathcal{T}=(T(t))_{t\geq 0}$ is such that $T(\tau)$ is bounded from below for some $\tau>0$ (and thus for all $\tau>0$), then the problem of similarity of $\mathcal{T}$ to a contraction semigroup reduces essentially to the same problem for a single operator. In this case, $\mathcal{T}\in\mathcal{SC}(H)$ if and only if there exists $\tau>0$ such that $T(\tau)$ is similar to a contraction (Theorem 4.11). As a consequence, we obtain a variant of Liapunov’s theorem (Corollary 4.12) for left invertible $C_{0}$-semigroups (so also for $C_{0}$-groups) $\mathcal{T}$ saying that for any $a$ strictly larger than the exponential type $\omega_{0}(\mathcal{T})$ of $\mathcal{T}$, there exists an equivalent Hilbertian norm $\|\cdot\|_{\mathscr{H}}$ on $H$ such that $\|T(t)\|_{\mathscr{H}}\leq e^{at}$ for all $t\geq 0.$ This result improves, in particular, [36, Corollary 7.2.5] and compare it with [3, Theorem 4.1].
@@
The proof of Theorem 1.2 is given in Section 4.3. In particular, it depends on Proposition 4.13, a partial case of a new characterization of tensor products of semigroups similar to semigroups of contractions obtained in [63, Theorem 1.1]. Note a Banach space version of Theorem 1.2 given in [7, Theorem 5.4], where Banach space semigroups bounded from below were characterized in terms of existence of equivalent norm making their negative generators dissipative.
@@ 2. Preliminaries: basic objects, results, and tools
See [71, Proposition 04], [44, p. 230], and [9, pp. 235-236] for these basic properties. Note that, since $\|\cdot\|_{\rm eq}$ is an equivalent norm, we have for all $h\in H$,
@@
The similarity constants were studied in particular in [44], where they are called distortion coefficients, and in [5]. Note that if $C(T)$ is finite, then it is attained, that is, there exists a Hilbertian norm $\|\cdot\|_{\rm{eq}}$ on $H$ such that $\|h\|\leq\|h\|_{\rm{eq}}\leq C(T)\|h\|$ and $\|Th\|_{\rm{eq}}\leq\|h\|_{\rm{eq}}$ for all $h\in H$, see [44, Proposition 2.4].
@@
If $\mathcal{T}$ is not jointly similar to contractions, then we set $\mathcal{C}(\mathcal{T})=\infty$. Similarly to the case of a single operator, $\mathcal{C}(\mathcal{T})$ is always attained. This can be inferred from the proof of [44, Proposition 2.4].
@@
There’s a geometric description of dilations due to Sarason, which in our setting takes the next form. Suppose $\mathcal{T}=(T(t))_{t\geq 0}$ is a semigroup on a Hilbert space $K$, and $H\subseteq K$ is a subspace. Then the compression $\mathcal{T}_{H}=P_{H}\mathcal{T}\mathord{\upharpoonright}_{H}$ is a semigroup on $H$ if and only if there exist $\mathcal{T}$-invariant subspaces $H_{2}$ and $H_{1}$ of $K$ such that the orthogonal difference $H_{2}\ominus H_{1}$ is precisely $H$. See e.g. [71, Theorem 1.7, Remark 1.8] for the proof of a more general result and comments. Such a subspace $H$ is often called semi-invariant in the literature. Clearly, $\mathcal{T}_{H}$ is strongly continuous if $\mathcal{T}$ is so.
@@
A detailed formal argument explaining how (2.5) fits into this definition can be found e.g. in [25, Section D.8].
@@
The class of $C_{0}$-semigroups on $H$ similar to a quasi-contraction semigroup will be denoted by $\mathcal{SQC}(H).$ Clearly, a semigroup $\mathcal{T}$ is in $\mathcal{SQC}(H)$ if and only if there exists an equivalent Hilbertian norm on $H$ making $\mathcal{T}$ a quasi-contraction semigroup. The class $\mathcal{SQC}(H)$ is quite large, and it includes e.g. $C_{0}$-groups on $H$ as noted in Section 1. Apart from the perturbation argument mentioned there, this follows also from the fact that the generator of a $C_{0}$-group on $H$ admits a bounded $H^{\infty}$-calculus in an appropriate strip. See [36, Section 7.2] for more on that, or [80, Section 8] for an alternative approach relevant for this paper.
Note that both contractivity and quasi-contractivity are highly unstable with respect to similarities. It is instructive to note that, as proved in [58, Theorem 2], for every $C_{0}$-semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ with unbounded generator on $H$ there is a semigroup on $H$ similar to $\mathcal{T},$ which is not quasi-contractive.
@@
where the convergence is uniform in $t$ for compact subsets of $(0,\infty)$; see, for instance, [26, Corollary III.5.5]. This formula will allow us to transfer similarity properties from resolvents to semigroups.

### v2

@@ 1.1. State of the art, motivations and illustrations
Despite Pisier’s negative result, the study of operators similar to contractions and their sub-classes became a well-developed area, and it continued to bear fruit, see e.g. [5], [16], [22], [33], [60] and [74] as samples. It is instructive to observe that if $T\in\mathcal{L}(H)$ satisfies the so-called Ritt’s resolvent condition $\|(z-T)^{-1}\|\leq C|z-1|^{-1}$ for $z$ outside the unit disc, then the polynomial boundedness of $T$ does imply its similarity to a contraction, [50, Theorem 5.1]. However, there are Ritt operators which are not polynomially bounded, and thus not similar to a contraction, as for instance [50, Proposition 5.2] shows. Among other results linking similarity to contraction to resolvent estimates, polynomial and power boundedness, we note [11], [16], [22] and [32]. However, it is worth noting that polynomial boundedness is a rather strong assumption, and it is often difficult to verify in concrete applications.
@@
Unfortunately, despite its importance for applications, $\mathcal{SC}(H)$ is not well-understood and in contrast to the discrete case there are very few conditions ensuring similarity to a contraction semigroup. At the same time, several pertinent examples were found illustrating the difficulty of the problem. Relying on ideas of Foguel [31], Packel constructed in [66] an example of a bounded $C_{0}$-semigroup $\mathcal{T}\not\in\mathcal{SC}(H),$ with the generator being rather implicit. In [9], Benchimol provided a similar construction of $\mathcal{T}\not\in\mathcal{SC}(H).$ However, he first constructed the generator of $\mathcal{T}$ as a bounded perturbation of a generator of a contraction semigroup, thus making both the generator and the semigroup transparent and explicit. Moreover, among other instructive examples, Chernoff showed in [17, p. 254] how to produce an exponentially stable $C_{0}$-semigroup that does not belong to $\mathcal{SC}(H).$ This somewhat discouraged research on $\mathcal{SC}(H),$ and much later even stronger counterexamples emerged. In particular, Le Merdy observed in [50, Theorem 1.1] that a similar example can be realised for exponentially stable holomorphic semigroup. In addition, in contrast to the case of holomorphic semigroups, one can find $\mathcal{T}\not\in\mathcal{SC}(H)$ whose negative generator is invertible and has bounded imaginary powers. The strongest known example of $\mathcal{T}\not\in\mathcal{SC}(H)$ was again found by Le Merdy in [51], see [37, Section 9.1] for further discussion. In his example, $\mathcal{T}=(T(t))_{t\geq 0}$ is exponentially stable, consists of compact operators for $t>0,$ and is holomorphic of angle $\pi/2.$ (The latter was not stated in [51], but can be easily deduced using e.g. [37, Lemma 9.1.2].) Very recently, among other improvements to Le Merdy’s example, it was shown in [64, Theorem 1.3] that one can construct an even stronger example by replacing exponential decay with nilpotency while preserving compactness, though necessarily omitting holomorphicity. However, holomorphicity is present in another example from [64, Theorem 1.3], where nilpotency is relaxed to quasi-nilpotency.
@@
Among very few positive results, Le Merdy proved in [50, Theorem 1.1] that a sectorially bounded holomorphic semigroup belongs to $\mathcal{SC}(H)$ if and only if its negative generator has bounded imaginary powers, or equivalently admits a bounded $H^{\infty}$-calculus on an appropriate sector. This allowed one to formulate similarity criteria in terms of square function estimates, [37, Theorem 7.3.1], [50, Theorem 4.3], and, a posteriori, invoking resolvent of the generator [13] (where boundedness of imaginary powers was described via integral resolvent bounds). Moreover, aiming at an analogue of Nagy’s theorem, it was shown in [80] that every bounded, uniformly continuous and quasi-compact semigroup in a Hilbert space is similar to a contraction semigroup. Furthermore, [3, Theorem 4.3] showed that the assumption of uniform continuity in [80] can be relaxed to the weaker conditions of analyticity and quasi-contractivity.
@@
The examples by Chernoff and Le Merdy, along with their improvements in [64, Theorem 1.3], demonstrate two key points. First, they show that continuous analogs of the classical criteria by Nagy and Rota fail in a dramatic way. Second, they clarify that our study concerns in fact the joint similarity of the semigroup $(T(t))_{t\geq 0}$ to a family of contractions on $H$, i.e. $\|RT(t)R^{-1}\|\leq 1$ for an invertible $R\in\mathcal{L}(H)$ and $t\geq 0$. This joint similarity does not, in general, follow from the similarity of each individual $T(t),$ $t\geq 0,$ to contraction - unlike the discrete case of the semigroup $(T^{n})_{n\geq 0}.$ As it will be made clear in Section 6 below, the property $\mathcal{T}\in\mathcal{SC}(H)$ depends on the behavior of the similarity constants $C(T(t))$ near zero. For an operator $T\in\mathcal{L}(H)$ its similarity constant is defined as
@@
Quasi-contraction semigroups originate from the classical work [56] and frequently appear in applications. For instance, they play a crucial role in proving the well-posedness of concrete linear PDEs via the (shifted) Lumer-Phillips theorem. They also arise in various stability problems, such as those related to Chernoff’s formula and its applications in numerical analysis, as well as in the study of nonlinear PDEs and their linearizations. To illustrate the concept of quasi-contractivity, observe that every $C_{0}$-semigroup $\mathcal{T}$ with a bounded generator $E$ is quasi-contractive as it satisfies the bound $\|T(t)\|\leq e^{\|E\|t}$ for all $t\geq 0.$ Furthermore, since any bounded perturbation of the generator of a quasi-contraction semigroup results in another quasi-contraction semigroup, bounded perturbations of normal operators whose spectrum lies in some left half-plane generate quasi-contraction semigroups. By a similar argument, one can show that every $C_{0}$-group on $H$ is similar to a quasi-contraction semigroup (cf. [37, Theorem 7.2.8 and Corollary 7.2.9]).
Thus, quasi-contractivity is a significantly weaker requirement than contractivity and naturally arises in semigroup theory. On the other hand, Chernoff’s example mentioned above ([17, p. 254]) actually yields a semigroup not similar to a quasi-contraction semigroup. Furthermore, by [37, Theorem 7.3.13] one can construct an analogous example where, in addition, negative generator is invertible and has bounded imaginary powers. Moreover, as we show in Section 7, the Packel and Le Merdy semigroups do not belong to $\mathcal{SQC}(H)$. Even stronger examples of semigroups outside $\mathcal{SQC}(H)$ can be found in [64]. At the same time, Benchimol’s semigroup is similar to a quasi-contraction semigroup since its generator arises as a bounded perturbation of the generator of a contraction semigroup. Thus the relations between $\mathcal{SC}(H)$ and $\mathcal{SQC}(H)$ appear to be rather intricate, and the aim to clarify them underlines the specifics and novelty of our approach. The paper links quasi-contractivity to the study of semigroups in $\mathcal{SC}(H)$ and sheds new light on their structure, revealing the different roles of small and large times and their interplay. It aims to emphasize these roles and to make the subject accessible to different audiences.
@@ 1.2. Ideas and results
The result can be proved by following the arguments from the proof of its discrete analogue, established in [43]. See also [72, Proposition 4.2] for a more general setting. This characterization, together with its generalization in the spirit of (1.1), is fundamental to our approach. Accordingly, we provide both results with new, more explicit proofs, placing them within the context of representations of (algebraic) semigroups. In particular, we make quotient norms explicit and thereby avoid the indirect argument requiring verification of the parallelogram law, as in [44]. In this way, given $\tau>0$, we infer that $(T(t))_{t\geq\tau}$ is jointly similar to contractions if (1.3) holds only for $t\geq\tau$. This observation becomes crucial in the next step of our program, where we describe $\mathcal{T}=(T(t))_{t\geq 0}$ with at least one $T(t)$ similar to a contraction.
@@
Remark that if $\mathcal{T}=(T(t))_{t\geq 0}$ is such that $T(\tau)$ is bounded from below for some $\tau>0$ (and thus for all $\tau>0$), then the problem of similarity of $\mathcal{T}$ to a contraction semigroup reduces essentially to the same problem for a single operator. In this case, $\mathcal{T}\in\mathcal{SC}(H)$ if and only if there exists $\tau>0$ such that $T(\tau)$ is similar to a contraction (Theorem 4.11). As a consequence, we obtain a variant of Liapunov’s theorem (Corollary 4.12) for left invertible $C_{0}$-semigroups (so also for $C_{0}$-groups) $\mathcal{T}$ saying that for any $a$ strictly larger than the exponential type $\omega_{0}(\mathcal{T})$ of $\mathcal{T}$, there exists an equivalent Hilbertian norm $\|\cdot\|_{\mathscr{H}}$ on $H$ such that $\|T(t)\|_{\mathscr{H}}\leq e^{at}$ for all $t\geq 0.$ This result improves, in particular, [37, Corollary 7.2.5] and compare it with [3, Theorem 4.1].
@@
The proof of Theorem 1.2 is given in Section 4.3. In particular, it depends on Proposition 4.13, a partial case of a new characterization of tensor products of semigroups similar to semigroups of contractions obtained in [64, Theorem 1.1]. Note a Banach space version of Theorem 1.2 given in [7, Theorem 5.4], where Banach space semigroups bounded from below were characterized in terms of existence of equivalent norm making their negative generators dissipative.
@@ 2. Preliminaries: basic objects, results, and tools
See [72, Proposition 04], [45, p. 230], and [9, pp. 235-236] for these basic properties. Note that, since $\|\cdot\|_{\rm eq}$ is an equivalent norm, we have for all $h\in H$,
@@
The similarity constants were studied in particular in [45], where they are called distortion coefficients, and in [5]. Note that if $C(T)$ is finite, then it is attained, that is, there exists a Hilbertian norm $\|\cdot\|_{\rm{eq}}$ on $H$ such that $\|h\|\leq\|h\|_{\rm{eq}}\leq C(T)\|h\|$ and $\|Th\|_{\rm{eq}}\leq\|h\|_{\rm{eq}}$ for all $h\in H$, see [45, Proposition 2.4].
@@
If $\mathcal{T}$ is not jointly similar to contractions, then we set $\mathcal{C}(\mathcal{T})=\infty$. Similarly to the case of a single operator, $\mathcal{C}(\mathcal{T})$ is always attained. This can be inferred from the proof of [45, Proposition 2.4].
@@
There’s a geometric description of dilations due to Sarason, which in our setting takes the next form. Suppose $\mathcal{T}=(T(t))_{t\geq 0}$ is a semigroup on a Hilbert space $K$, and $H\subseteq K$ is a subspace. Then the compression $\mathcal{T}_{H}=P_{H}\mathcal{T}\mathord{\upharpoonright}_{H}$ is a semigroup on $H$ if and only if there exist $\mathcal{T}$-invariant subspaces $H_{2}$ and $H_{1}$ of $K$ such that the orthogonal difference $H_{2}\ominus H_{1}$ is precisely $H$. See e.g. [72, Theorem 1.7, Remark 1.8] for the proof of a more general result and comments. Such a subspace $H$ is often called semi-invariant in the literature. Clearly, $\mathcal{T}_{H}$ is strongly continuous if $\mathcal{T}$ is so.
@@
A detailed formal argument explaining how (2.5) fits into this definition can be found e.g. in [26, Section D.8].
@@
The class of $C_{0}$-semigroups on $H$ similar to a quasi-contraction semigroup will be denoted by $\mathcal{SQC}(H).$ Clearly, a semigroup $\mathcal{T}$ is in $\mathcal{SQC}(H)$ if and only if there exists an equivalent Hilbertian norm on $H$ making $\mathcal{T}$ a quasi-contraction semigroup. The class $\mathcal{SQC}(H)$ is quite large, and it includes e.g. $C_{0}$-groups on $H$ as noted in Section 1. Apart from the perturbation argument mentioned there, this follows also from the fact that the generator of a $C_{0}$-group on $H$ admits a bounded $H^{\infty}$-calculus in an appropriate strip. See [37, Section 7.2] for more on that, or [81, Section 8] for an alternative approach relevant for this paper.
Note that both contractivity and quasi-contractivity are highly unstable with respect to similarities. It is instructive to note that, as proved in [59, Theorem 2], for every $C_{0}$-semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ with unbounded generator on $H$ there is a semigroup on $H$ similar to $\mathcal{T},$ which is not quasi-contractive.
@@
where the convergence is uniform in $t$ for compact subsets of $(0,\infty)$; see, for instance, [27, Corollary III.5.5]. This formula will allow us to transfer similarity properties from resolvents to semigroups.